# D Scaling Up Deepfake Detection by Learning from Discrepancy

### v1

@@ Appendix 0.B Different Classifier Heads
Appendix 0.C Comparative Analysis of Uniquely Detected Samples
Appendix 0.D Detailed Mean Accuracy Results of Comparisons with the State-of-the-arts
In this section, we report the detailed mean accuracy results of comparisons with the state-of-the-art in Tab. 6, as a supplement to Table 1 in Sec 4.5, including ADM [13], BigGAN [4], GLIDE [31], Midjourney [2], LDM [36], VQDM [20], wukong [3], ProGAN [23], CycleGAN [51], StyleGAN [24], StyleGAN2 [25], GauGAN [33], StarGAN [8], Deepfakes [37], whichfaceisreal [1], SITD [6], SAN [11], CRN [7], IMLE [28], and DALLE [35]. Results of generators with the same architecture but different parameters are averaged. For example, the result of BigGAN [4] in this table is the average of BigGAN in UFD [32] and BigGAN in GenImage [52].

### v2

@@ Appendix B Different Classifier Heads
Appendix C Comparative Analysis of Uniquely Detected Samples
Appendix D Detailed Mean Accuracy Results of Comparisons with the State-of-the-arts
In this section, we report the detailed mean accuracy results of comparisons with the state-of-the-art in Sec. 3, as a supplement to Table 1 in Sec 4.5, including ADM [13], BigGAN [4], GLIDE [29], Midjourney [2], LDM [34], VQDM [18], wukong [3], ProGAN [21], CycleGAN [49], StyleGAN [22], StyleGAN2 [23], GauGAN [31], StarGAN [8], Deepfakes [35], whichfaceisreal [1], SITD [6], SAN [11], CRN [7], IMLE [26], and DALLE [33]. Results of generators with the same architecture but different parameters are averaged. For example, the result of BigGAN [4] in this table is the average of BigGAN in UFD [30] and BigGAN in GenImage [50].
Appendix E Visualization of the Detector’s Attention
effectively learns more universal artifacts through discrepancies introduced by image transformations, achieving superior performance. We employ the occlusion technique [46] to visually demonstrate this to identify regions of interest during inference, as illustrated in Fig. 2. Our analysis reveals that UFD focuses on limited regions, often leading to misjudgments due to its narrow attention. In contrast, exhibits a significantly broader attention range, further validating that the introduced discrepancy enables D3 to capture more comprehensive and universal artifacts.